OpenStudy (anonymous):

6 − |3x + 9| = 3 solve plz

OpenStudy (anonymous):

First add -6 to both sides

OpenStudy (anonymous):

So, what do u get?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

-I3x+9I=-3 I3x+9I=3 Now, 3x+9=3 or 3x+9=-3

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

but isnt it suppose to be no solution since the absolute value most be equal to a possive number? in this case we have a -3

OpenStudy (anonymous):

ooo i got it, i see now

OpenStudy (theviper):

6 − |3x + 9| = 3 6 - (3x+9)=3 6-3x-9=3 -3x-3=3 -3x=3+3 -3x=9 x=9/-3=-3

OpenStudy (anonymous):

so xis equal to 9 and -3?